# Composite likelihood estimation of stationary Gaussian processes

### v1

@@ 2.3 Asymptotic theory
Assumption 2.2, together with the moment condition $E(|\log f(y;\theta)|)<\infty$, are sufficient to ensure a unique maximum at $\theta_{0}$ by the information inequality, see Newey and McFadden (1994, Lemma 2.2). The existence of the moment holds trivially for the Gaussian distribution. Moreover, in our setting the assumption can be verified if $K$ and $q$ are sufficiently large relative to the dimension of $\theta$, $p$. Of course, this assumes that all the parameters of the model enter into the densities of the chosen events in $Q$, and that they do so in a way that is linearly independent.
@@
Assume that the true model is the stationary Gaussian process defined in (2.1) and that Assumptions 2.12.2 hold. Then, as $n\rightarrow\infty$,
@@
Suppose the conditions from Theorem 2.1 hold and that $\theta_{0}$ is an interior point, i.e. $\theta_{0}\in\mathrm{int}(\Theta)$. Let $L_{\infty}$ be a slowly varying function at infinity, i.e. $\lim_{x\rightarrow\infty}L_{\infty}(tx)\div L_{\infty}(x)=1$, for all $t>0$. Then, as $n\rightarrow\infty$, it holds that [^7] The notation $f(h)\sim g(h)$ means asymptotic equivalence, i.e. $f(h)/g(h)\rightarrow 1$ as $|h|\rightarrow\infty$.
@@
    If $\int_{0}^{\infty}\gamma_{h}\mathrm{d}h<\infty$ or if $\gamma_{h}\sim h^{-\beta}L_{\infty}(h)$ as $h\rightarrow\infty$ for $\beta\in(1/2,1)$, then
[…]
@@
    $H(\theta_{0})=-\sum_{k\in Q}\mathbb{E}\left(\frac{\partial^{2}}{\partial\theta^{\top}\partial\theta}\log\left(f_{k}(Y_{k};\theta)\right)\mid_{\theta=\theta_{0}}\right),$
@@
    $V(\theta_{0})$ $=\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}\left(s_{n}(\theta)s_{n}(\theta)^{\top}\right)$
@@
    ×$\times\left(\left[\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k^{1}}^{-1}(\theta)\big{)}_{j_{3},j_{4}}\right]_{r=1}^{p}\right)^{\top}\times\left(\gamma_{l+j_{1}-j_{3}}\gamma_{l+j_{2}-j_{4}}+\gamma_{l+j_{1}-j_{4}}\gamma_{l+j_{2}-j_{3}}\right),$
    where $s_{n}(\theta)=\frac{\partial}{\partial\theta}cl(\theta;y)$. Furthermore, the infinite series defining $V(\theta_{0})$ is convergent.
@@
[…]
    with $G(\theta_{0})$ and $H(\theta_{0})$ defined as above, but where the variability matrix is redefined as
    $V(\theta_{0})$ $=\lim_{n\rightarrow\infty}\frac{1}{nL_{\gamma}(n)}\mathbb{E}\left(s_{n}(\theta)s_{n}(\theta)^{\top}\right)$
    $=\lim_{n\rightarrow\infty}\frac{1}{4}\frac{1}{L_{\gamma}(n)}\sum_{k^{1},k^{2}\in Q}\sum_{l=-n}^{n}\sum_{j_{1},j_{2}\in k_{1}}\sum_{j_{3},j_{4}\in k_{2}}\left[\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k^{1}}^{-1}(\theta)\big{)}_{j_{1},j_{2}}\right]_{r=1}^{p}$
    ×$\times\left(\left[\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k^{2}}^{-1}(\theta)\big{)}_{j_{3},j_{4}}\right]_{r=1}^{p}\right)^{\top}\times\left(\gamma_{l+j_{1}-j_{3}}\gamma_{l+j_{2}-j_{4}}+\gamma_{l+j_{1}-j_{4}}\gamma_{l+j_{2}-j_{3}}\right),$
    and $L_{\gamma}$ is a slowly varying function, related to $L_{\infty}$, as defined in Appendix A.
@@
[…]
    where $L_{2}(n)=2C_{2}L^{2}_{\infty}(n)$, $C_{2}=[(1-2\beta)(2-\beta)]^{-1}$, $\Psi$ is a $p\times 1$ vector with elements
    $\Psi_{r}=\frac{1}{2}\sum_{k^{j}\in Q}\sum_{j_{1},j_{2}\in k^{j}}\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k}^{-1}(\theta)\big{)}_{j_{1},j_{2}},$ (2.6)
    for $r=1,\dots,p$ and $Z_{2,H}$ is a Rosenblatt process with parameter $21-\beta/$ 2.
The first part covers the short memory setting with an integrable acf, but it also permits long memory (as defined in (3.2)) with $\beta\in(1/2,1)$. Here, we get a standard central limit theorem with a standard $n^{-1/2}$ rate of convergence. However, because the composite likelihood misspecifies the true likelihood, the second Bartlett identity (or information matrix equality) is violated, so the Fisher information is replaced with the Godambe (1960) information, which has a sandwich form. The ratio of these matrices determines the loss of efficiency of $\hat{\theta}_{\text{MCLE}}$ relative to $\hat{\theta}_{\text{MLE}}$, which achieves the Cramer-Rao lower bound asymptotically. This is the price we pay for not doing full maximum likelihood.
The second part is the borderline case with $\beta=1/2$, where the polynomial decay in the long memory setting starts to impair the convergence toward the asymptotic distribution of $\hat{\theta}_{\mathrm{MCLE}}$. Here, the variability matrix has to be scaled by an additional term, $L_{\gamma}(n)$, whereby the overall rate of convergence is reduced, but the limit remains Gaussian. An inspection of the expression for $L_{\gamma}(n)$ in Appendix A shows that, even if $\lim_{h\rightarrow\infty}L_{\infty}(h)$ exists, the convergence rate of $\hat{\theta}_{\text{MCLE}}$ deteriorates to $\log(n)n^{-1/2}$, as consistent with Davis and Yau (2011).
In the last part, $\beta\in(0,1/2)$, so the acf subsides exceedingly slow. Here, we get a non-central limit theorem with a non-standard convergence rate, which is nevertheless a classical result in the realm of long memory processes.
In practice, the variance-covariance matrix of the composite score function in the limiting distribution, $V(\theta_{0})$, of $\hat{\theta}_{\text{MCLE}}$ is challenging to compute, since the gradient vanishes at $\theta=\hat{\theta}_{\text{MCLE}}$. Instead, we propose to estimate $G^{-1}(\theta_{0})$ via a parametric bootstrap.
@@
The composite likelihood estimator is derived from the score of the log-composite likelihood function, and since we work with Gaussian data, the stochastic terms in the score relate to the autocovariance function of the process. Hence, the main difference between these settings corresponds to the limit theory for autocovariances calculated with a known mean or estimated mean. This was analyzed in the Gaussian setting by Hosking (1996), who shows that if the memory is not too strong (i.e. short memory or long memory with $\beta>1/2$), there is no impact. That is, estimating the mean does not alter neither the rate of convergence nor the limiting distribution.
On the other hand, with pervasive long memory (i.e. $\beta<1/2$), estimation of the mean makes a difference relative to our case 3. While the convergence rate is unchanged, the limiting distribution is no longer a Rosenblatt distribution, but it has a more complicated expression. [^8] To get a better understanding of the difference, one can compare the cumulants of the limiting distribution in Theorem 4.1 of Hosking (1996) with those of the Rosenblatt distribution, which are given by $\kappa_{1}=0$, $\kappa_{2}=1$, and $\kappa_{k}=2^{k-1}(k-1)![\sigma(\beta)]^{k}c_{k}$, for $k\geq 2$, where $\sigma(\beta)=2^{-1/2}[(1-2\beta)(1-\beta)]^{1/2}$ and $c_{k}=\int_{0}^{1}\int_{0}^{1}\cdots\int_{0}^{1}|x_{1}-x_{2}|^{-\beta}|x_{2}-x_{3}|^{-\beta}\cdots|x_{k-1}-x_{k}|^{-\beta}|x_{k}-x_{1}|^{-\beta}\mathrm{d}x_{1}\mathrm{d}x_{2}\cdots\mathrm{d}x_{k}$. It also instills an important finite sample downward bias in the sample acf, which impairs the estimation in practice. We shed light on this in the simulation study.
@@ 3 A view toward stochastic volatility
In this section, we narrow down the estimation of continuous-time stationary Gaussian processes to a particular pair of parametric models. The first is the fractional Ornstein-Uhlenbeck process (fOU), which has attracted a lot of attention in the recent literature on roughness in the stochastic volatility of financial asset returns (e.g., Gatheral, Jaisson, and Rosenbaum, 2018; Fukasawa, Takabatake, and Westphal, 2022; Bolko, Christensen, Pakkanen, and Veliyev, 2023; Wang, Xiao, and Yu, 2023). The second is from a so-called Cauchy class, which was proposed in Gneiting and Schlather (2004). In financial econometrics, it has—to the best of our knowledge—only been studied as a volatility model by Bennedsen, Lunde, and Pakkanen (2022). [^9] The Cauchy class has found widespread application in other fields, such as modeling infectious disease spread (Meyer and Held, 2014), network traffic (Li and Lim, 2008), or forecasting wind speed (Liu, Song, and Zio, 2021).
@@
A stationary stochastic process has roughness index $\alpha$ if its acf has the following behaviour around the origin:
$(1-\rho(h))\sim L(h)|h|^{2\alpha+1},\quad|h|\rightarrow 0,$ (3.1)
for a function $L(h)$ that is slowly varying at zero and $\alpha\in(-1/2,1/2)$.
A Brownian motion has $\alpha=0$. Moreover, because it is a continuous Gaussian process, there is a one-to-one correspondence between the roughness index and the Hölder continuity of its sample paths (Bennedsen, 2020). Thus, we refer to a process as being rough if it has a negative roughness index, $\alpha<0$, as its sample paths are then more irregular—i.e. less Hölder continuous—than those of a Brownian motion. Conversely, we call a process smooth if it has a positive roughness index, $\alpha>0$.
@@
$\rho(h)\sim L_{\infty}(h)|h|^{-\beta},\quad|h|\rightarrow\infty,$ (3.2)
@@ 3.1 The fOU process
$\mathrm{d}Y_{t}=-\kappa(Y_{t}-\mu)\mathrm{d}t+\nu\mathrm{d}B_{t}^{H},$ (3.3)
where $B^{H}$ is a fractional Brownian motion (fBm). [^10] A fBm is a centered (mean zero) stationary Gaussian process, which is uniquely characterized by the covariance function $\mathbb{E}\big{[}B_{t}^{H}B_{s}^{H}\big{]}=\frac{1}{2}\big{(}t^{2H}+s^{2H}-|t-s|^{2H}\big{)}$. The single parameter $H\in(0,1)$ is called the Hurst exponent.
The unique stationary solution to this stochastic differential equation was derived in Cheridito, Kawaguchi, and Maejima (2003): [^11] The fBm is not a semimartingale, except for $H=1/2$ where it collapses to a standard Brownian motion (Rogers, 1997). Therefore, the solution in (3.4) cannot be defined as a standard Itô integral, but it should be interpreted in a stricter sense as a pathwise Riemann-Stieltjes integral. Although this is more restrictive, it is meaningful here, since the fBm is continuous and the exponential function is of bounded variation.
@@
where $X_{t}=\sqrt{\dfrac{\kappa^{2H}}{H\Gamma(2H)}}\int_{-\infty}^{t}e^{-\kappa(t-s)}\mathrm{d}B_{s}^{H}$ and $\Gamma$ is the Gamma function. The additional scaling in front of the stochastic integral in (3.4) relative to (3.3) is a reparameterization of the model to ensure $\nu$ can be interpreted as the standard deviation of $Y_{t}$. It follows from $\text{var}\left(\int_{-\infty}^{t}e^{-\kappa(t-s)}\mathrm{d}B_{s}^{H}\right)=\kappa^{-2H}H\Gamma(2H)$.
@@
The acf of the fBm decays hyperbolically with $\rho_{h}=O(h^{2(H-1)})$ for $H\neq 1/2$, whereas for $H=1/2$ it is geometrically bounded. Moreover, it follows from Cheridito, Kawaguchi, and Maejima (2003) that the acf of the fOU process inherits the order of decay of the background driving fBm.
By the Kolmogorov-Chentsov theorem (Chentsov, 1956), fBm has a modification, where the sample paths are Hölder continuous with exponent $\gamma\in(0,H)$. So the Hurst parameter determine how regular the paths are and it is linked to the roughness index in (3.1) through the relation $\alpha=H-1/2$. Hence, for $H<1/2$ the fBm process is rough, whereas for $H>1/2$ it is smooth. Furthermore, it readily follows that it has long memory for $H>1/2$, whereas it is possesses short memory for $H\leq 1/2$. As before, these properties are passed down to the fOU process.
This bottom line is that the Hurst exponent controls both the roughness and memory properties of the fOU process. This renders it capable of exhibiting either roughness or long memory, but not both. This is also rather intuitive, since an examination of the acf of the time-changed process $B_{ct}^{H}$, for some $c>0$, reveals that the fBm is self-similar.

### v2

@@ 2.3 Asymptotic theory
Assumption 2.2 is a standard identification condition formulated in terms of a density. Together with the moment condition $\mathbb{E}(|\log f(y;\theta)|)<\infty$, it is sufficient to ensure that the log-composite likelihood has a unique maximum at $\theta_{0}$ by the information inequality, see Newey and McFadden (1994, Lemma 2.2). The existence of the moment holds trivially for the Gaussian distribution. Moreover, in our setting the identification condition can be verified if $K$ and $q$ are sufficiently large and diverse relative to the dimension of $\theta$, $p$. This requires that all the parameters of the model enter into the densities of the chosen events in $Q$, and they should do so in a way that is linearly independent.
In the Gaussian setting, this amounts to establish that the selected autocovariances identify $\theta_{0}$. This is the content of the next lemma.
Lemma 2.1.
Let $\gamma_{h}(\theta)$ denote the parameterized ACF of $Y$. Suppose that
$\gamma_{k_{i}^{j}}(\theta)=\gamma_{k_{i}^{j}}(\theta_{0})\text{ for }i=1,\dots,q^{j},\quad j=1,\dots,K\Longrightarrow\theta=\theta_{0}.$
Then, $\mathbb{E}[cl(\theta;Y)]$ has a unique maximum at $\theta=\theta_{0}$.
The main message of this lemma—as explained further in Appendix A—is that to ensure identification we need to select more lags, either through one long tuple or multiple shorter tuples, than the maximum number of times $\gamma_{h}(\theta)$ can cross for distinct $\theta$. Since we examine processes with non-oscillatory ACFs, the condition is typically fulfilled even with a relatively primitive implementation.
@@
Assume that the true model is the stationary Gaussian process defined in (2.1) and that Assumptions 2.12.2 hold. Then, under standard regularity conditions, as $n\rightarrow\infty$,
@@
Suppose the conditions from Theorem 2.1 hold and that $\theta_{0}$ is an interior point, i.e. $\theta_{0}\in\mathrm{int}(\Theta)$. Let $s_{n}(\theta)=\frac{\partial}{\partial\theta}cl(\theta;y)$ be the score and $L_{\infty}$ a slowly varying function at infinity, i.e. $\lim_{x\rightarrow\infty}L_{\infty}(tx)\div L_{\infty}(x)=1$, for all $t>0$. Then, as $n\rightarrow\infty$, it holds that [^10] The notation $f(h)\sim g(h)$ means asymptotic equivalence, i.e. $f(h)/g(h)\rightarrow 1$ as $h\rightarrow\infty$.
@@
    If $\int_{0}^{\infty}|\gamma_{h}|\mathrm{d}h<\infty$ or if $\gamma_{h}\sim h^{-\beta}L_{\infty}(h)$ as $h\rightarrow\infty$ for $\beta\in(1/2,1]$, then
    $\sqrt{n}\big(\hat{\theta}_{\mathrm{MCLE}}-\theta_{0}\big)\overset{d}{\longrightarrow}N\left(0,G(\theta_{0})^{-1}\right),$
@@
    $H(\theta)=-\sum_{k\in Q}\mathbb{E}\left(\frac{\partial^{2}}{\partial\theta\partial\theta^{\top}}\log\left(f_{k}(Y_{k};\theta)\right)\right),$
@@
    $V(\theta)$ $=\lim_{n\rightarrow\infty}\frac{1}{n}\mathbb{E}\left(s_{n}(\theta)s_{n}(\theta)^{\top}\right)$
@@
    $\times\left(\left[\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k^{2}}^{-1}(\theta)\big)_{j_{3},j_{4}}\right]_{r=1}^{p}\right)^{\top}$
    $\times\left(\gamma_{(l+j_{1}-j_{3})\Delta}\gamma_{(l+j_{2}-j_{4})\Delta}+\gamma_{(l+j_{1}-j_{4})\Delta}\gamma_{(l+j_{2}-j_{3})\Delta}\right).$
@@
    $\frac{\sqrt{n}}{\sqrt{L_{\gamma}(n)}}\big(\hat{\theta}_{\mathrm{MCLE}}-\theta_{0}\big)\overset{d}{\longrightarrow}H(\theta_{0})^{-1}\Psi(\theta_{0})N(0,1),$
    with $H(\theta_{0})$ defined as above, and where $\Psi$ is a $p\times 1$ vector with elements
    $\Psi_{r}(\theta)=\sum_{k^{j}\in Q}\sum_{j_{1},j_{2}\in k^{j}}\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k}^{-1}(\theta)\big)_{j_{1},j_{2}},$ (2.5)
    for $r=1,\dots,p$, and $L_{\gamma}$ is a slowly varying function, related to $L_{\infty}$, as defined in Appendix A.
@@
    $\frac{n^{\beta}}{\sqrt{L_{2}(n)}}(\hat{\theta}_{\mathrm{MCLE}}-\theta_{0})\overset{d}{\longrightarrow}H(\theta_{0})^{-1}\Psi(\theta_{0})Z_{2,H}(1),$ (2.6)
    where $L_{2}(n)=4C_{2}L_{\infty}^{2}(n)$, $C_{2}=[(1-2\beta)(2-\beta)]^{-1}$, $\Psi$ is as above, and $Z_{2,H}$ is a Rosenblatt process with parameter $H=1-\beta/2$.
The first part of Theorem 2.2 covers the short memory setting with an integrable ACF, but it also permits long memory for $\beta\in(1/2,1]$ with a square-integrable ACF. Here, we get a standard central limit theorem with a standard $n^{-1/2}$ rate of convergence. However, because the composite likelihood misspecifies the true likelihood, the second Bartlett identity (or information matrix equality) is violated, so the Fisher information is replaced with the Godambe (1960) information, which has the sandwich form. The ratio of these matrices determines the loss of efficiency of $\hat{\theta}_{\text{MCLE}}$ relative to $\hat{\theta}_{\text{MLE}}$ in this setting, since the latter achieves the Cramer-Rao lower bound asymptotically.
The second part is the borderline case with $\beta=1/2$, where the polynomial decay in the long memory setting starts to impair the properties of $\hat{\theta}_{\mathrm{MCLE}}$. First, the ACF is not square-integrable, so the variability matrix has to be scaled by an additional term, $L_{\gamma}(n)$, whereby the rate of convergence is reduced. An inspection of the expression for $L_{\gamma}(n)$ in Appendix A shows that, even if $\lim_{h\rightarrow\infty}L_{\infty}(h)$ exists, the convergence rate of $\hat{\theta}_{\text{MCLE}}$ deteriorates to $\log(n)n^{-1/2}$, as consistent with Davis and Yau (2011). Second, although the limit remains Gaussian, it is important to observe that the entire parameter vector is loading on a single standard normal variate, so the asymptotic distribution is degenerate. This follows from Theorem 4 in Hosking (1996), because the difference between the sample ACF at any distinct lags converge to zero in probability for $\beta\leq 1/2$.
In the last part, $\beta\in(0,1/2)$, the ACF subsides exceedingly slow. Here, we get a non-central limit theorem with a non-standard convergence rate, which is nevertheless a classical result in the realm of long memory processes. The asymptotic distribution remains degenerate, but it is of Rosenblatt form.
The bottom line is that for the MCLE, the asymptotic theory presents a discontinuity in both the rate of convergence and the form of the limiting distribution, such that inference is complicated. This is the price we pay for not doing MLE. On the other hand, as shown in Figure 1, estimation can be done in a split second, and it may be the only feasible likelihood-based procedure when looking at very large sample sizes, as demonstrated in one of the applications in our empirical investigation.
@@
The composite likelihood estimator is derived from the score of the log-composite likelihood function, and since we work with Gaussian data, the stochastic terms in the score relate to the autocovariance function of the process. Hence, the main difference between these settings corresponds to the limit theory for autocovariances calculated with a known mean or estimated mean. This was analyzed in the Gaussian setting by Hosking (1996), who shows that if the memory is not too strong (i.e. short memory or long memory with $\beta\geq 1/2$), there is no impact. That is, estimating the mean does not alter neither the rate of convergence nor the limiting distribution.
On the other hand, with pervasive long memory (i.e. $\beta<1/2$), estimation of the mean makes a difference relative to our case 3. While the convergence rate is unchanged, the limiting distribution is no longer a Rosenblatt distribution, but it has a more complicated expression. [^11] To get a better understanding of the difference, one can compare the cumulants of the limiting distribution in Theorem 4.1 of Hosking (1996) with those of the Rosenblatt distribution, which are given by $\kappa_{1}=0$, $\kappa_{2}=1$, and $\kappa_{k}=2^{k-1}(k-1)![\sigma(\beta)]^{k}c_{k}$, for $k\geq 2$, where $\sigma(\beta)=2^{-1/2}[(1-2\beta)(1-\beta)]^{1/2}$ and $c_{k}=\int_{0}^{1}\int_{0}^{1}\cdots\int_{0}^{1}|x_{1}-x_{2}|^{-\beta}|x_{2}-x_{3}|^{-\beta}\cdots|x_{k-1}-x_{k}|^{-\beta}|x_{k}-x_{1}|^{-\beta}\mathrm{d}x_{1}\mathrm{d}x_{2}\cdots\mathrm{d}x_{k}$.
Mean estimation also instills a finite sample downward bias in the sample ACF, which impairs the estimation in practice. We shed light on this in the simulation study.
Remark 2.
The assumption of Gaussianity for the background driving stationary process can be weakened by assuming that $X_{t}$ is strictly stationary and ergodic with a finite second moment. We can then interpret (2.4) as a Gaussian quasi-composite likelihood function. In that case, $\hat{\theta}_{\text{MCLE}}$ is consistent for the subset of the parameter vector that determines the mean and covariance function of $X_{t}$, since the Gaussian log-likelihood is uniquely maximized at the true value of these (Bollerslev and Wooldridge, 1992, Lemma A.2). Furthermore, the rates of convergence in Theorem 2.2 remain unchanged. We can also establish asymptotic normality in the short-memory case by imposing that $X_{t}$ is strong mixing with an appropriate rate of decay on the mixing coefficients together with a moment condition (see, e.g., Keenan, 2008). In the long-memory setting, however, deriving an explicit expression for the limiting distribution appears to be a more onerous problem.
@@ 3 A view toward stochastic volatility
In this section, we narrow down the estimation of continuous-time stationary Gaussian processes to a particular pair of parametric models. The first is the fractional Ornstein-Uhlenbeck process (fOU), which has attracted a lot of attention in the recent literature on roughness in the stochastic volatility of financial asset returns (e.g., Gatheral, Jaisson, and Rosenbaum, 2018; Fukasawa, Takabatake, and Westphal, 2022; Bolko, Christensen, Pakkanen, and Veliyev, 2023; Wang, Xiao, and Yu, 2023; Shi, Yu, and Zhang, 2024; Wang, Xiao, Yu, and Zhang, 2025). The second is from a so-called Cauchy class, which was proposed in Gneiting and Schlather (2004). In financial econometrics, it has—to the best of our knowledge—only been studied as a volatility model by Bennedsen, Lunde, and Pakkanen (2022). [^12] The Cauchy class has seen widespread application in other fields, such as modeling infectious disease spread, network traffic, or forecasting wind speed.
@@
A stationary stochastic process has roughness index $\alpha$ if its ACF has the following behavior around the origin:
$(1-\rho_{h})\sim L(h)|h|^{2\alpha+1},\quad|h|\rightarrow 0,$ (3.1)
for a function, $L(h)$, that is slowly varying at zero, and $\alpha\in(-1/2,1/2)$.
The concept of roughness is closely related to the behavior of the (second-order) variogram of a zero-mean real-valued stochastic process with stationary increments that was studied in Gatheral, Jaisson, and Rosenbaum (2018). For example, (the increments of) a Brownian motion has $\alpha=0$. Moreover, because it is a continuous Gaussian process, there is a one-to-one correspondence between the roughness index and the Hölder continuity of its sample paths. Thus, we refer to a process as being rough if it has a negative roughness index, $\alpha<0$, as its sample paths are then more irregular—i.e. less Hölder continuous—than those of a Brownian motion. Conversely, we call a process smooth if it has a positive roughness index, $\alpha>0$.
@@
$\rho_{h}\sim L_{\infty}(h)|h|^{-\beta},\quad|h|\rightarrow\infty,$ (3.2)
@@ 3.1 The fOU process
$\mathrm{d}Y_{t}=-\kappa(Y_{t}-\mu)\mathrm{d}t+\tilde{\nu}\mathrm{d}B_{t}^{H},$ (3.3)
where $B^{H}$ is a fractional Brownian motion (fBm). [^13] A fBm is a centered Gaussian process, which is uniquely characterized by its covariance function $\mathbb{E}\big[B_{t}^{H}B_{s}^{H}\big]=\frac{1}{2}\big(t^{2H}+s^{2H}-|t-s|^{2H}\big)$. The parameter $H\in(0,1)$ is the Hurst exponent. The fBm has stationary increments called fractional Gaussian noise (fGn).
The unique stationary solution to this stochastic differential equation was derived in Cheridito, Kawaguchi, and Maejima (2003): [^14] The fBm is not a semimartingale, except for $H=1/2$ where it collapses to a sBm. Therefore, the solution in (3.4) cannot be defined as a standard Itô integral, but it should be interpreted as a pathwise Riemann-Stieltjes integral. Although this is more restrictive, it is meaningful here, since the fBm is continuous and the exponential function is of bounded variation.
@@
where $X_{t}=b\int_{-\infty}^{t}e^{-\kappa(t-s)}\mathrm{d}B_{s}^{H}$, $\nu=\tilde{\nu}/b$, $b=\sqrt{\dfrac{\kappa^{2H}}{H\Gamma(2H)}}$, and $\Gamma$ is the Gamma function. The additional scaling in front of the stochastic integral in (3.4) relative to (3.3) is a reparameterization of the model to ensure that $\nu$ is the standard deviation of $Y_{t}$. It follows from $\text{var}\left(\int_{-\infty}^{t}e^{-\kappa(t-s)}\mathrm{d}B_{s}^{H}\right)=\kappa^{-2H}H\Gamma(2H)$.
@@
The ACF of the fBm decays hyperbolically with $\rho_{h}=O(h^{2(H-1)})$ for $H\neq 1/2$, whereas for $H=1/2$ it is geometrically bounded. It follows from the Kolmogorov-Chentsov theorem that fBm has a modification, where the sample paths are Hölder continuous with exponent $\gamma\in(0,H)$, so $H$ determines the regularity of the process. In particular, the Hurst parameter is linked to the roughness index in Definition 1 through the relation $\alpha=H-1/2$. Hence, for $H<1/2$ the fBm is rough, whereas for $H>1/2$ it is smooth. Furthermore, it is readily seen that its connection with Definition 2 is $\beta=2(1-H)$. This implies that the fBm has long memory for $H>1/2$, whereas it possesses short memory for $H<1/2$. As explained in Cheridito, Kawaguchi, and Maejima (2003), the ACF of the fOU process inherits the decay rate of the background driving fBm, so these properties are passed on one-to-one.
The bottom line is that the Hurst exponent controls both the roughness and memory properties of the fOU process. This renders it capable of exhibiting either roughness or long memory, but not both. This is also rather intuitive, since an examination of the ACF of the time-changed process $B_{ct}^{H}$, for some $c>0$, reveals that the fBm is self-similar.